# Leveraging Large Language Models for Relevance Judgments in Legal Case Retrieval

### v2

@@ Abstract.
ccs: Information systems Specialized information retrieval
@@ 3.2. Automated Relevance Judgments for Legal Cases
[…]
@@ 4.2. Implementation Details
According to the crime classification defined in the chapters of the Criminal Law’s specific provisions, the demonstration set in ADM is constructed based on 40 query cases, which cover the main types of cases present in the LeCaRd dataset. For each query case, four associated cases with different gold-standard relevance labels are provided, along with FE and FA instructions corresponding to two factual aspects: MF and LF.
For the relevance evaluation, we compare the judgments of the proposed method with the golden labels of the top 30 candidate cases in every query case. For the LLM, we opted for GPT-3.5-0613 with temperature 0.4. In the heatmap experiment, we randomly sample 400 human-labeled pairs for all relevance levels, from 0 to 3.
@@ 4.3. Evaluation of the Relevance Judgments (RQ1 & 2)
Separately based on two aspects, Material Facts (MF) and Legal Facts (LF), we assess the reliability of the proposed method by calculating Cohen’s Kappa for multiple judgments conducted at a certain temperature in Figure 9. To demonstrate the validity, we evaluate the automated judgment consistency by calculating Kappa values with the golden labels in Figure 10, and then draw heatmaps that visualize the consistency between the proposed method and human labeling.
For the reliability test shown in Figure 9, as the temperature increases, there is a moderate change in the Kappa values among our multiple relevance judgments, indicating good reliability and that our carefully designed expert instruction can effectively orient the output of LLM. For the validity test shown in Figure 10, it can be observed that the consistency between Legal Facts and human labels is highly significant, while Material Facts exhibit relative significance when compared to human labels. One primary reason for this is that each case involves one or more defendants; when comparing the facts related to the defendants between two cases, there are one-to-many and many-to-many comparative relationships, which could involve more subjective judgment. During our research, we do identify a significant degree of subjective judgment in the human judgments of Material Facts under such comparisons. This is attributed to the fact that human labels do not originate from a single individual, and there is inevitably some bias among different individuals’ judgments. Moreover, this bias is more likely to manifest in Material Fact, due to the arbitrary forms of writing it. However, according to the labeling standard, the relevance of legal facts (score=2) is more critical than that of material facts (score=1). Therefore, the quality of judgments for Legal Facts still predominantly influences the final judgment results. This indicates that our proposed automatic relevance judgment can closely approximate human labeling. Interestingly, in all temperature settings, the Avg. kappa values remain relatively stable. This illustrates the reliability of aligning with human labels again.
@@
Figure 9. The Cohen’s Kappa between multiple times of MF&LF relevance judgment over different temperature settings.
Figure 10. The Cohen’s Kappa between the MF&LF and the golden labels over different temperature settings: Avg. represents the average of Kappa values between multiple relevance judgments and the golden labels
@@ 4.4.1. Baselines
In We select two basic deep models (BERT (Devlin et al., 2018), Longformer (Beltagy et al., 2020)) and an advanced SOTA method SAILER (Li et al., 2023b) as our baselines. Note that our goal is to evaluate the effectiveness of automatic relevance judgments. The focus is on validating whether the relative performance differences in downstream models before and after applying the augmented data, rather than achieving state-of-the-art results on the LCR task.
@@ 5. Limitation and additional experiments
Using only 9 expert-designed demonstration sets, we generated pairwise annotations from unlabeled cases in the training set and selected 20k samples to fine-tune BERT. As shown in Table4, results compared with reported SOTA(Feng et al., 2024; Bi et al., 2022) show that our method remains effective in different legal domains.
@@ 6. Conclusions and Future Works
7. Usage of Generative AI
In this project, Generative AI was utilized in a limited and transparent manner. Specifically, AI tools were employed for the following purposes:
Language polishing: We used AI assistance to improve the clarity, grammar, and fluency of written text, ensuring professional and readable documentation.
Programming support: During the coding process, we made use of Cursor, an AI-powered code editor, to assist with code suggestions, debugging, and refactoring.
No AI tools were used for generating core research content, analysis, or decision-making processes. All conceptual and methodological work was conducted independently by the authors.

### v3

@@ Abstract.
$c$s: Information systems Specialized information retrieval
@@ 3.2. Automated Relevance Judgments for Legal Cases
$\mathcal{D}_{\text{FE}}^{\text{LF}}$ $=\{\text{Expert demonstrations for Legal Facts in Fact Extraction}\},$
$\mathcal{D}_{\text{FE}}^{\text{MF}}$ $=\{\text{Expert demonstrations for Material Facts in Fact Extraction}\},$
$\mathcal{D}_{\text{FA}}^{\text{LF}}$ $=\{\text{Expert demonstrations for Legal Facts in Fact Annotation}\},$
$\mathcal{D}_{\text{FA}}^{\text{MF}}$ $=\{\text{Expert demonstrations for Material Facts in Fact Annotation}\}.$
@@ 4.2. Implementation Details
According to the crime classification defined in the chapters of the Criminal Law’s specific provisions, the expert-curated demonstration set for ADM is constructed based on 40 query cases, which cover the main types of cases present in the LeCaRd dataset. For each query case, four associated cases with different gold-standard relevance labels are provided, along with FE and FA instructions corresponding to two factual aspects: MF and LF.
For the relevance evaluation, we compare the judgments of the proposed method with the golden labels of the top 30 candidate cases in every query case. For the LLM, we opted for GPT-3.5-turbo with temperature 0.4. In the heatmap experiment, we randomly sample 400 human-labeled pairs for all relevance levels, from 0 to 3. While more capable models (e.g., gpt-4o) were available in later stages of this research, we retained gpt-3.5-turbo for all experiments to maintain comparability with earlier results. Given that gpt-4o consistently surpasses gpt-3.5-turbo in reasoning and generation quality, we expect our method’s performance to further improve when paired with stronger backbones, which will be investigated in future studies.
@@ 4.3. Evaluation of the Relevance Judgments (RQ1 & 2)
Separately based on two aspects, Material Facts (MF) and Legal Facts (LF), we assess the reliability of the proposed method by calculating Cohen’s Kappa for multiple judgments conducted at a certain temperature in Figure 9. To demonstrate the validity, we evaluate the automated judgment consistency by calculating Kappa values with the golden labels in Figure 10, and then draw heatmaps that visualize the consistency between the proposed method and human labeling. To more intuitively demonstrate the reliability and validity of our method, we also report relevance judgment results obtained directly using a Chain-of-Thought (CoT) prompting strategy(with all necessary labeling guidance and few shot demonstrations).
For the validity test in Figure 10, the consistency between Legal Facts and human labels is highly significant. This demonstrates that our automatic relevance judgment closely approximates human experts’ labeling. We also observe that Material Facts show only relative significance. A key reason is that each case involves one or more defendants; when comparing defendant-related facts between cases, one-to-many and many-to-many relationships often require subjective judgment. We observed such subjectivity in human annotations of Material Facts, partly because labels come from multiple annotators with inevitable individual biases. Moreover, this bias is more likely to manifest in Material Fact, due to the arbitrary forms of writing it. Nonetheless, according to the labeling standard of LeCaRD, Legal Facts (score=2) are more critical than Material Facts (score=1), so the overall quality of Legal Fact judgments dominates final results. Notably, directly employing a CoT prompting strategy for case relevance judgment yields poor performance in both reliability and validity, underscoring the necessity of our approach.
@@
Figure 9. Reliability Test: The Cohen’s Kappa between multiple times of MF&LF relevance judgment over different temperature settings.
Figure 10. Validity Test: The Cohen’s Kappa between the MF&LF and the golden labels over different temperature settings: Avg. represents the average of Kappa values between multiple relevance judgments and the golden labels
@@ 4.4.1. Baselines
In data augmentation experiments, we select two basic deep models (BERT (Devlin et al., 2018), Longformer (Beltagy et al., 2020)) and an advanced SOTA method SAILER (Li et al., 2023b) as our baselines. Note that our primary objective is to assess the quality of the automatic relevance judgments. We evaluate this by measuring the relative performance gains in downstream models after applying the augmented data, rather than aiming to achieve a SOTA model on the LCR task.
@@ 5. Limitation and additional experiments
Using only 9 expert-designed demonstration sets, we generated pairwise annotations from unlabeled cases in the training set and selected 20k samples to fine-tune BERT. As shown in Table4, results compared with the reported SOTA result(Feng et al., 2024; Bi et al., 2022) show that our method remains effective in different legal domains.